An asset used in a 4-year project falls in the 5-year MACRS class for tax purposes. The asset has an acquisition cost of $7,900,000 and will be sold for $1,640,000 at the end of the project. If the tax rate is 23 percent, what is the aftertax salvage value of the asset? Refer to (MACRS schedule)
Answer #1

Acquisition Cost of Asset = $7,900,000

Depreciation charges for 4 years under MACRS Schedule = 20%+32%+19.20%+11.52%

Hence, Written down value of asset = 100 - 82.72% = 17.28%

Written Down Value of Asset at the end of the project = 7,900,000*17.28%

= $1,365,120

Sale Value of Asset = $1,640,000

Gain on Sale = $274,880

Tax @23% = 63,222.4

After Tax Salvage Value of Asset = $1,640,000 - 63,222.4

= $1,576,777.6